# A New Look at the Quantum Measurement Problem

### v2

@@ 6 Experimental verification
The results presented in Figs 1 and 2 demonstrate that the electron originally located on the 1s orbital, with its most probable radius at $r=a_{o}$, is shifted to its most probable radius at $r=2\ a_{o}$ as a result of the measurements. This means that the electron interacting with a measuring apparatus gets confined into a measurement eigenstate. Comparison of the original electron’s probability density ${\cal{P}}_{1s}(r_{a})$ on the orbital $1s$ (see Fig. 1) to ${\mathfrak{P}}(r_{a})$ plotted in Fig. 2 shows that as a result of the measurement the electron is confined to a well-defined position at $r=2a_{o}$, and that its probability density is very narrow and sharply centered at this position, which is consistent with Born’s position measurement principles (e.g., [33,34]). The time scale for the existence of this measurement eigenstate is given by Eq. (7), which represents the duration of this eigenstate.
The above theoretical results can be verified experimentally by using a quantum microscope similar to that designed by an international team of researchers [38], who used it to measure the orbital structure of Stark states in an excited hydrogen atom (see their Fig. 3). They reported a similar trend in shifting the maxima and broadening the density distributions for shorter wavelengths as those presented in Fig. 2. There are several other measuring methods developed to study the quantum-classical correspondence [39-43], but some of these methods may not be suitable to observe single orbitals in a hydrogen atom [42].
@@ 7 Conclusions
In this paper, a new asymmetric equation, which is complementary to the Schrödinger equation, is used to investigate the measurement problem of quantum mechanics. The obtained results demonstrate that while the Schrödinger equation describes the evolution of the wavefunction prior to any measurement, the new asymmetric equation represents the behavior of the wavefunction during the measurement process. The temporal solutions to the new asymmetric equation give a time-scale for the duration of the measurement eigenstate, whose probability density is determined from the spatial solutions.
To describe the transition from unitary (reversible) representation of the electron on its $1s$ orbital, which is given by the Schrödinger equation, to its non-unitary (irreversible) evolution resulting from the measurement, which is given by the new asymmetric equation, the product of both probabilty densities is calculated. The main results are that the time-scales are or the order of $10^{-17}$ $10^{-15}$ s, depending on the frequency of the EM waves used in the measurement, and that the computed probability densities are sharply centered around $r=2\ a_{o}$, which is consitent with Born’s position measurement postulates.
The predicted radial probability densities resulting from the measurement process can be verified experimentally by using a quantum microscope (e.g., [38]). Based on the presented results, it is proposed that while the unitary evolution of the wavefunction is described by the Schrödinger equation, its non-unitary behavior caused by measurements is represented by the new asymmetric equation; this means that both equations are required in nonrelativistic quantum mechanics to fully describe the behavior of the wavefunction.
Acknowledgment: The author thanks Dora Musielak for valuable comments on the earlier version of this manuscript.

### v3

@@ 6 Experimental verification
The results presented in Figs 1 and 2 demonstrate that the electron originally located on the 1s orbital, with its most probable radius at $r=a_{o}$, is shifted to its most probable radius at $r=2\ a_{o}$ as a result of the measurements. This means that the electron interacting with a measuring apparatus gets confined into an eigenstate that corresponds to the measurement. Comparison of the original electron’s probability density ${\cal{P}}_{1s}(r_{a})$ on the orbital $1s$ (see Fig. 1) to ${\mathfrak{P}}(r_{a})$ plotted in Fig. 2 shows that as a result of the measurement the electron is confined to a well-defined position at $r=2a_{o}$, and that its probability density is very narrow and sharply centered at this position, which is consistent with Born’s position measurement principles (e.g., [41,42]). The time scale for the existence of this eigenstate is given by Eq. (9), which represents its duration.
The theoretical results obtained in this paper can be verified experimentally by using a quantum microscope similar to that designed by an international team of researchers [54], who used it to measure the orbital structure of Stark states in an excited hydrogen atom (see their Fig. 3). They reported a similar trend in shifting the maxima and broadening the density distributions for shorter wavelengths as those shown in Fig. 2. There are several other measuring methods developed to study the quantum-classical correspondence [54-59], but some of these methods may not be suitable to observe single orbitals in a hydrogen atom [58].
After using the NAE to solve the quantum measurement problem in this paper, and quantum jumps in [47], the NAE has the potential to develop new quantum based technologies, which would verify its validity and applicability. As described in a recent comprehensive review [60], cryptographic systems and secure direct communications as well as emerging fields like the quantum internet are all based on modern quantum technologies. However, to explore potentials of developing new quantum technologies based on the NAE is out the scope of this paper.
@@ 7 Conclusions
In this paper, a new asymmetric equation, which is complementary to the Schrödinger equation, is used to investigate the measurement problem of quantum mechanics. The obtained results demonstrate that while the Schrödinger equation describes the evolution of the wavefunction prior to any measurement, the new asymmetric equation represents the behavior of the wavefunction during the measurement process. The temporal solutions to the new asymmetric equation give a time-scale for the duration of the eigenstate that corresponds to the measurement; probability density at this eigenstate is determined from the spatial solutions.
To describe the transition from unitary (reversible) representation of the electron on its $1s$ orbital, which is given by the Schrödinger equation, to its non-unitary (irreversible) evolution resulting from the measurement, which is given by the new asymmetric equation, the product of both probabilty densities is calculated. The main results are that the time-scales are or the order of $10^{-17}$ $10^{-15}$ s, depending on the frequency of the EM waves used in the measurement, and that the computed probability densities are sharply centered around $r=2\ a_{o}$, which shows that Born’s position measurement postulates naturally emerge from the presented theoretical results. In other words, the new asymmetric equation and first principles used to derive it also account for Born’s rules of quantum mechanics.
The predicted radial probability densities resulting from the measurement process can be verified experimentally by using a quantum microscope (e.g., [54]). Another way to verify the validity and applicability of the new asymmetric equation would require exploring its potentials in contributing to already known quantum based technologies (e.g., [60]), or to be used to develop new quantum technologies based on this equation; however, this is out of the scope of this paper.
Based on the presented results, it is proposed that while the unitary evolution of the wavefunction is described by the Schrödinger equation, its non-unitary behavior caused by measurements is represented by the new asymmetric equation; this means that both equations are required in nonrelativistic quantum mechanics to fully describe the spatial and temporal evolution of the wavefunction, and the quantum measurement problem.
Acknowledgments I appreciate very much constructive comments and insightful suggestions made by three anonymous referees, which allowed me to significantly improve this paper. The author also thanks Dora Musielak for reading the earlier version of this manuscript and suggesting improvements in its structure and presentation of the results.